OpenStudy (anonymous):

The area of a circular sun spot is growing at a rate of 600 km2/sec. (a) How fast is the radius growing at the instant when it equals 4,000 km?

ganeshie8 (ganeshie8):

\(\huge A = \pi r^2 \\ \huge \frac{dA}{dt} = 2\pi r * \frac{dr}{dt} = 600 \\ \huge => \frac{dr}{dt} = \frac{600}{2\pi r}\)

ganeshie8 (ganeshie8):

to know how fast radius is changing, plugin r = 4000 in above equation

OpenStudy (anonymous):

awesome thank you!

ganeshie8 (ganeshie8):

np :)